# GreenHyperSpectra: A multi-source hyperspectral dataset for global vegetation trait prediction

### v2

@@ Abstract
Plant traits such as leaf carbon content and leaf mass are essential variables in the study of biodiversity and climate change. However, conventional field sampling cannot feasibly cover trait variation at ecologically meaningful spatial scales. Machine learning represents a valuable solution for plant trait prediction across ecosystems, leveraging hyperspectral data from remote sensing. Nevertheless, trait prediction from hyperspectral data is challenged by label scarcity and substantial domain shifts (e.g. across sensors, ecological distributions), requiring robust cross-domain methods. Here, we present GreenHyperSpectra, a pretraining dataset encompassing real-world cross-sensor and cross-ecosystem samples designed to benchmark trait prediction with semi- and self-supervised methods. We adopt an evaluation framework encompassing in-distribution and out-of-distribution scenarios. We successfully leverage GreenHyperSpectra to pretrain label-efficient multi-output regression models that outperform the state-of-the-art supervised baseline. Our empirical analyses demonstrate substantial improvements in learning spectral representations for trait prediction, establishing a comprehensive methodological framework to catalyze research at the intersection of representation learning and plant functional traits assessment. We also share the code and data for this study here.
@@ 1 Introduction
Plant functional traits are a fundamental component of biodiversity assessment, offering insights into plant productivity, ecological interactions, resilience, and adaptation to environmental change (cavender2020remote, ; funk2017revisiting, ; skidmore2021priority, ; yan2023essential, ). Leaf traits such as leaf mass per area, as well as chlorophyll, nitrogen, and carbon content, are key to understanding plant growth dynamics and ecosystem processes such as carbon cycling and productivity (berger2022multi, ; bongers2021functional, ; damm2018remote, ; roscher2012using, ; zarco2018previsual, ; zarco2019chlorophyll, ). The monitoring of these traits is thus crucial for understanding ecosystem function and guiding biodiversity conservation strategies (cavender2022integrating, ; pettorelli2021time, ; xu2021ensuring, ).
@@
Initiatives such as the Intergovernmental Science-Policy Platform on Biodiversity and Ecosystem Services (IPBES; (diaz2015ipbes, ; ipbes2019global, )) have raised global awareness around the urgent need for monitoring functional traits and their diversity across spatial scales. However, we still lack efficient tools to track these functional traits in space and time. Hyperspectral remote sensing from airborne and satellite systems has emerged as a promising tool to bridge this gap, enabling non-destructive, scalable, and repeatable reflectance measurements that can be used to predict these traits (berger2020crop, ; cherif2023spectra, ; jetz2016monitoring, ; serbin2019arctic, ; van2018functional, ). Hyperspectral sensors measure radiation reflected from the ground across hundreds of narrow, contiguous spectral bands, spanning the visible to shortwave infrared (VNIR + SWIR) domains. These measurements are informative for trait prediction, as they are directly influenced by the chemical and structural characteristics of plant leaves and canopies (jacquemoud2019leaf, ). Plant trait prediction from hyperspectral data is inherently a regression problem, and was initially often explored with Partial Least Squares Regression (PLSR, (geladi1986partial, )) to link hyperspectral observations to individual traits (feilhauer2010brightness, ; serbin2015remotely, ; helsen2021evaluating, ; ustin2009retrieval, ). However, non-parametric machine learning methods, in particular deep learning, have recently been explored to offer greater flexibility in modeling complex, non-linear trait-spectral relationships and trait-trait interactions (cherif2023spectra, ; tsakiridis2020simultaneous, ; he2016novel, ). In this context, trait prediction is framed as a multi-output regression task within a multi-task learning framework, where the outputs, representing multiple plant traits, are inherently correlated.
@@
Trait prediction poses fundamental machine learning challenges including heterogeneous target distributions requiring specialized multi-task methods cherif2023spectra , extreme label scarcity, and significant distributional biases (e.g. spatial, Figure 2). Current supervised approaches demonstrate limited cross-domain generalization due to training on sparse, non-representative datasets (danilevicz2022plant, ; pichler2020machine, ). Hyperspectral data further compound these challenges exhibiting substantial covariate shifts across acquisition conditions, sensor configurations and resolutions, radiometric calibrations and variable input modalities. To address these limitations, we introduce GreenHyperSpectra, a large spectral dataset designed to improve representation robustness against domain adaptation challenges while being collected from multiple ecosystems, instruments, spatial resolutions, and acquisition conditions. This dataset enables semi- and self-supervised learning applications, which take advantage of vast unlabeled spectral data, providing extensive coverage and variability to facilitate benchmarking.
@@ 2 Related work
Despite increasing hyperspectral data availability, plant trait prediction is still largely constrained by the lack of large annotated datasets (danilevicz2022plant, ; todman2023small, ). Trait labels are costly and time-consuming to obtain, often requiring field sampling and laboratory analysis  (cornelissen2003handbook, ; baraloto2010functional, ). Available datasets lack harmonization and often differ in sampling strategy, measurement assumptions and protocols. As a result, most labeled datasets are geographically and ecologically limited, with sparse coverage across space and time, as well as across ecosystems and acquisition conditions. To address this, previous studies explored synthetic datasets generated from Radiative Transfer Models (RTMs, (feret2008prospect, )) that simulate canopy spectral responses under diverse conditions. In this context, it motivated hybrid approaches that combine RTM-based simulations with machine learning (danner2021efficient, ; mederer2025plant, ; parigi2024towards, ; tagliabue2022hybrid, ; verrelst2015optical, ; wang2021airborne, ).
@@
However, several comparisons reveal important limitations, with models trained on real, multi-site datasets consistently outperforming those using only synthetic spectra (mederer2025plant, ; parigi2024towards, ), highlighting a persistent domain gap between simulated and field data.
@@
In this regard, there is a need of large-scale unlabeled spectral datasets from real measurements to pretrain models. To support this, a number of benchmarking efforts have been introduced (fuchs2023hyspecnet, ; wang2025hypersigma, ; braham2024spectralearth, ). However, most of these initiatives are built from a single sensor type, restricting spectral diversity and limiting generalization to new acquisition conditions. Moreover, these datasets often consist of full hyperspectral imagery, where spatially contiguous pixels are subject to spatial autocorrelation. This spatial redundancy reduces the spectral variability necessary for training models that generalize well in trait prediction task, which depend on spectral rather than spatial information (thoreau2024toulouse, ). They also typically include a broad mix of land cover types, including non-vegetated surfaces. This introduces inefficiencies, as vegetated pixels must be sampled through additional preprocessing.
@@
Large-scale unlabeled datasets have been leveraged for self-supervised learning to improve natural image representations Chen_2021_ICCV ; Assran2023Self ; oquab2024dinov as well as in remote sensing with multispectral data tseng2023lightweight ; NEURIPS2023_11822e84 ; xiong2024neuralplasticityinspiredmultimodalfoundation ; szwarcman2025prithvieo20versatilemultitemporalfoundation ; bountos_fomo_2025 ; astruc2025anysatearthobservationmodel ; tseng2025galileolearninggloballocal . Semi- and self-supervised learning techniques are increasingly being explored to exploit hyperspectral data for image classification, segmentation, and super-resolution (ahmad2021hyperspectral, ; ranjan2023unlocking, ; wang2022self, ). These methods are designed to better learn the spectral representation of hyperspectral remote sensing data to reduce the need for reference labels. Approaches such as masked autoencoders (MAE) (cong2022satmae, ; hong2021spectralformer, ; reed2023scale, ; scheibenreif2023masked, ; thoreau2024toulouse, ), contrastive learning frameworks (chen2020simple, ; fuller2023croma, ; he2020momentum, ), generative networks (GAN) (alipour2020structure, ; he2017generative, ; kwak2023semi, ; roy2021generative, ; zhan2018semi, ; zhu2018generative, ), and autoencoders (AE) (ahmad2019segmented, ; gallo2023self, ; thoreau2024toulouse, ; zhao2017spectral, ) have been successfully applied for land cover classification. Whereas most attempts at trait prediction using hyperspectral data relied on fully supervised pipelines, notably PLSR (serbin2015remotely, ; singh2015imaging, ; ustin2009retrieval, ; wang2020foliar, ), Gaussian Process Regression (GPR) (tagliabue2022hybrid, ; verrelst2013gaussian, ; verrelst2012gaussian, ), ANN (schlerf2006inversion, ) and deep learning methods (cherif2023spectra, ; pullanagari2021field, ; shi2022convolution, ), they are cardinally constrained by label scarcity, hampering their ability to reliably generalize across ecosystems, sensor platforms, and acquisition conditions. Applying semi- and self-supervised methods to trait prediction remains largely unexplored and offers a compelling direction for investigation. Recent semi- and self-supervised applications in trait prediction include vision transformers for nitrogen estimation from simulated data (gallo2023self, ) and Long short-term memory (LSTM) models for chlorophyll prediction with limited spectral bands (zhao2023improving, ) (i.e. VNIR-only). While semi- and self-supervised methods show promise for trait prediction, existing models remain constrained to single traits. Moreover, existing models are typically limited to specific sensor configurations and experimental conditions, limiting generalization across sensor modalities (e.g. full-range vs. half-range spectrometers), acquisition geometries, and vegetation types. This underscores the need for flexible approaches that can handle heterogeneous inputs while supporting transferable predictions in diverse real-world ecological scenarios. To the best of our knowledge, no semi- or self-supervised method addresses trait prediction via multi-output regression.
@@ 3 The GreenHyperSpectra dataset
The multi-platform nature of our dataset introduces valuable reflectance signal variability through differences in spatial and spectral resolution, sun-sensor geometry, scene heterogeneity, background conditions and pre-processing from radiance to reflectance. This variability, often lacking in single-platform or synthetic datasets, is essential to develop generalizable models capable of scaling across diverse remote sensing contexts (cherif2023spectra, ). While satellite-based datasets such as SpectralEarth (braham2024spectralearth, ) provide temporally rich but sensor-specific imagery, our dataset uniquely incorporates multi-sensor observations across spaceborne, airborne, and proximal platforms.
@@ 4 Benchmarking methods and protocols
We consider a supervised CNN-based method (cherif2023spectra, ; mederer2025plant, ) as a baseline, selected for its state-of-the-art performance in multi-trait plant prediction from a sparse annotated dataset. It is built upon EfficientNet-B0 (tan2019efficientnet, ) specifically framed for 1D feature extraction. The network employs multi-output regression to simultaneously predict the seven plant traits, a strategy that demonstrably outperforms single-trait modeling approaches (cherif2023spectra, ; scutari2014multiple, ).
@@
We frame the SR-GAN framework (olmschenk2019generalizing, ) to address hyperspectral plant trait prediction. Our implementation employs a 1D convolutional GAN architecture designed specifically for spectral data processing. In this setup, the generator learns to produce synthetic reflectance spectra, while the discriminator simultaneously performs trait regression and learns discriminative feature representations. The training objective is formulated as a composite loss that encourages the discriminator to pull real spectral samples closer in the feature space, while pushing representations of generated (synthetic) samples further apart. This contrastive learning approach allows the model to leverage unlabeled data by learning informative spectral embeddings. The overall architecture of the SR-GAN framework is illustrated in Figure 3(a) and detailed formulations of the loss components are provided in Appendix B.1.
@@
We introduce a version of the autoencoder framework proposed by (she2024spectra, ), which replaces the decoder with a non-learnable RTM module to reconstruct spectra, thereby integrating physical constraints into the modeling process. Specifically, our implementation employs PROSAIL-PRO (feret2021prospect, ), constraining the latent space to correspond directly to plant traits. PROSAIL-PRO is an RTM that combines the leaf reflectance model (PROSPECT,(jacquemoud1990prospect, )) with the canopy reflectance model (4SAIL,(verhoef2007unified, )). PROSPECT simulates leaf reflectance and transmittance based on biochemical composition and internal structure, while 4SAIL models the propagation of light through a vegetation canopy. Together, they simulate canopy spectral reflectance in the 400–2500 nm range using inputs such as chlorophyll content, leaf area index, and leaf angle. As previously mentioned regarding the gap between RTM-simulated and real-world spectra (§ 2), we address the inherent discrepancies between RTM-generated and observed spectra, primarily resulting from simplified geometric assumptions within the model, by implementing a learnable correction layer that refines the simulated output (she2024spectra, ). Our enhanced framework introduces three key improvements over the original design (she2024spectra, ): (1) incorporation of PROSAIL-PRO, (2) application of a supervised loss component targeting trait predictions, and (3) implementation of a composite reconstruction loss combining cosine similarity and mean absolute error to capture both spectral shape characteristics and amplitude information. The overall architecture is illustrated in Figure 3(b) and the specifications are detailed in Appendix B.2 .
@@
We adopt a MAE framework, originally designed for land cover classification (thoreau2024toulouse, ), to predict plant traits with hyperspectral data. The model leverages self-supervised learning by reconstructing randomly masked spectral regions, enabling the extraction of meaningful representations from unlabeled hyperspectral signatures. Similarly to the RTM-AE, our adaptation incorporates a modified reconstruction objective that combines cosine similarity and mean squared error (MSE) with appropriate weighting, allowing the model to capture both spectral shape characteristics and amplitude information. For downstream trait prediction, we attach a multi-output regression head to the latent features and fine-tune the model using labeled data. The overall architecture is illustrated in Figure 3(c) and ablation studies for the MAE architecture are provided in Appendix B.3.

### v3

@@ Abstract
Plant traits such as leaf carbon content and leaf mass are essential variables in the study of biodiversity and climate change. However, conventional field sampling cannot feasibly cover trait variation at ecologically meaningful spatial scales. Machine learning represents a valuable solution for plant trait prediction across ecosystems, leveraging hyperspectral data from remote sensing. Nevertheless, trait prediction from hyperspectral data is challenged by label scarcity and substantial domain shifts (e.g. across sensors, ecological distributions), requiring robust cross-domain methods. Here, we present GreenHyperSpectra, a pretraining dataset encompassing real-world cross-sensor and cross-ecosystem samples designed to benchmark trait prediction with semi- and self-supervised methods. We adopt an evaluation framework encompassing in-distribution and out-of-distribution scenarios. We successfully leverage GreenHyperSpectra to pretrain label-efficient multi-output regression models that outperform the state-of-the-art supervised baseline. Our empirical analyses demonstrate substantial improvements in learning spectral representations for trait prediction, establishing a comprehensive methodological framework to catalyze research at the intersection of representation learning and plant functional traits assessment. We also share the dataset [^1] GreenHyperSpectra dataset: https://huggingface.co/datasets/Avatarr05/GreenHyperSpectra, code and pretrained model objects for this study here.
@@ 1 Introduction
Plant functional traits are a fundamental component of biodiversity assessment, offering insights into plant productivity, ecological interactions, resilience, and adaptation to environmental change (cavender2020remote; funk2017revisiting; skidmore2021priority; yan2023essential). Leaf traits such as leaf mass per area, as well as chlorophyll, nitrogen, and carbon content, are key to understanding plant growth dynamics and ecosystem processes such as carbon cycling and productivity (berger2022multi; bongers2021functional; damm2018remote; roscher2012using; zarco2018previsual; zarco2019chlorophyll). The monitoring of these traits is thus crucial for understanding ecosystem function and guiding biodiversity conservation strategies (cavender2022integrating; pettorelli2021time; xu2021ensuring).
@@
Initiatives such as the Intergovernmental Science-Policy Platform on Biodiversity and Ecosystem Services (IPBES; (diaz2015ipbes; ipbes2019global)) have raised global awareness around the urgent need for monitoring functional traits and their diversity across spatial scales. However, we still lack efficient tools to track these functional traits in space and time. Hyperspectral remote sensing from airborne and satellite systems has emerged as a promising tool to bridge this gap, enabling non-destructive, scalable, and repeatable reflectance measurements that can be used to predict these traits (berger2020crop; cherif2023spectra; jetz2016monitoring; serbin2019arctic; van2018functional). Hyperspectral sensors measure radiation reflected from the ground across hundreds of narrow, contiguous spectral bands, spanning the visible to shortwave infrared (VNIR + SWIR) domains. These measurements are informative for trait prediction, as they are directly influenced by the chemical and structural characteristics of plant leaves and canopies (jacquemoud2019leaf). Plant trait prediction from hyperspectral data is inherently a regression problem, and was initially often explored with Partial Least Squares Regression (PLSR, (geladi1986partial)) to link hyperspectral observations to individual traits (feilhauer2010brightness; serbin2015remotely; helsen2021evaluating; ustin2009retrieval). However, non-parametric machine learning methods, in particular deep learning, have recently been explored to offer greater flexibility in modeling complex, non-linear trait-spectral relationships and trait-trait interactions (cherif2023spectra; tsakiridis2020simultaneous; he2016novel). In this context, trait prediction is framed as a multi-output regression task within a multi-task learning framework, where the outputs, representing multiple plant traits, are inherently correlated.
@@
Trait prediction poses fundamental machine learning challenges including heterogeneous target distributions requiring specialized multi-task methods cherif2023spectra, extreme label scarcity, and significant distributional biases (e.g. spatial, Figure 2). Current supervised approaches demonstrate limited cross-domain generalization due to training on sparse, non-representative datasets (danilevicz2022plant; pichler2020machine). Hyperspectral data further compound these challenges exhibiting substantial covariate shifts across acquisition conditions, sensor configurations and resolutions, radiometric calibrations and variable input modalities. To address these limitations, we introduce GreenHyperSpectra, a large spectral dataset designed to improve representation robustness against domain adaptation challenges while being collected from multiple ecosystems, instruments, spatial resolutions, and acquisition conditions. This dataset enables semi- and self-supervised learning applications, which take advantage of vast unlabeled spectral data, providing extensive coverage and variability to facilitate benchmarking.
@@ 2 Related work
Despite increasing hyperspectral data availability, plant trait prediction is still largely constrained by the lack of large annotated datasets (danilevicz2022plant; todman2023small). Trait labels are costly and time-consuming to obtain, often requiring field sampling and laboratory analysis  (cornelissen2003handbook; baraloto2010functional). Available datasets lack harmonization and often differ in sampling strategy, measurement assumptions and protocols. As a result, most labeled datasets are geographically and ecologically limited, with sparse coverage across space and time, as well as across ecosystems and acquisition conditions. To address this, previous studies explored synthetic datasets generated from Radiative Transfer Models (RTMs, (feret2008prospect)) that simulate canopy spectral responses under diverse conditions. In this context, it motivated hybrid approaches that combine RTM-based simulations with machine learning (danner2021efficient; mederer2025plant; parigi2024towards; tagliabue2022hybrid; verrelst2015optical; wang2021airborne).
@@
However, several comparisons reveal important limitations, with models trained on real, multi-site datasets consistently outperforming those using only synthetic spectra (mederer2025plant; parigi2024towards), highlighting a persistent domain gap between simulated and field data.
@@
In this regard, there is a need of large-scale unlabeled spectral datasets from real measurements to pretrain models. To support this, a number of benchmarking efforts have been introduced (fuchs2023hyspecnet; wang2025hypersigma; braham2024spectralearth). However, most of these initiatives are built from a single sensor type, restricting spectral diversity and limiting generalization to new acquisition conditions. Moreover, these datasets often consist of full hyperspectral imagery, where spatially contiguous pixels are subject to spatial autocorrelation. This spatial redundancy reduces the spectral variability necessary for training models that generalize well in trait prediction task, which depend on spectral rather than spatial information (thoreau2024toulouse). They also typically include a broad mix of land cover types, including non-vegetated surfaces. This introduces inefficiencies, as vegetated pixels must be sampled through additional preprocessing.
@@
Large-scale unlabeled datasets have been leveraged for self-supervised learning to improve natural image representations Chen_2021_ICCV; Assran2023Self; oquab2024dinov as well as in remote sensing with multispectral data tseng2023lightweight; NEURIPS2023_11822e84; xiong2024neuralplasticityinspiredmultimodalfoundation; szwarcman2025prithvieo20versatilemultitemporalfoundation; bountos_fomo_2025; astruc2025anysatearthobservationmodel; tseng2025galileolearninggloballocal. Semi- and self-supervised learning techniques are increasingly being explored to exploit hyperspectral data for image classification, segmentation, and super-resolution (ahmad2021hyperspectral; ranjan2023unlocking; wang2022self). These methods are designed to better learn the spectral representation of hyperspectral remote sensing data to reduce the need for reference labels. Approaches such as masked autoencoders (MAE) (cong2022satmae; hong2021spectralformer; reed2023scale; scheibenreif2023masked; thoreau2024toulouse), contrastive learning frameworks (chen2020simple; fuller2023croma; he2020momentum), generative networks (GAN) (alipour2020structure; he2017generative; kwak2023semi; roy2021generative; zhan2018semi; zhu2018generative), and autoencoders (AE) (ahmad2019segmented; gallo2023self; thoreau2024toulouse; zhao2017spectral) have been successfully applied for land cover classification. Whereas most attempts at trait prediction using hyperspectral data relied on fully supervised pipelines, notably PLSR (serbin2015remotely; singh2015imaging; ustin2009retrieval; wang2020foliar), Gaussian Process Regression (GPR) (tagliabue2022hybrid; verrelst2013gaussian; verrelst2012gaussian), ANN (schlerf2006inversion) and deep learning methods (cherif2023spectra; pullanagari2021field; shi2022convolution), they are cardinally constrained by label scarcity, hampering their ability to reliably generalize across ecosystems, sensor platforms, and acquisition conditions. Applying semi- and self-supervised methods to trait prediction remains largely unexplored and offers a compelling direction for investigation. Recent semi- and self-supervised applications in trait prediction include vision transformers for nitrogen estimation from simulated data (gallo2023self) and Long short-term memory (LSTM) models for chlorophyll prediction with limited spectral bands (zhao2023improving) (i.e. VNIR-only). While semi- and self-supervised methods show promise for trait prediction, existing models remain constrained to single traits. Moreover, existing models are typically limited to specific sensor configurations and experimental conditions, limiting generalization across sensor modalities (e.g. full-range vs. half-range spectrometers), acquisition geometries, and vegetation types. This underscores the need for flexible approaches that can handle heterogeneous inputs while supporting transferable predictions in diverse real-world ecological scenarios. To the best of our knowledge, no semi- or self-supervised method addresses trait prediction via multi-output regression.
@@ 3 The GreenHyperSpectra dataset
The multi-platform nature of our dataset introduces valuable reflectance signal variability through differences in spatial and spectral resolution, sun-sensor geometry, scene heterogeneity, background conditions and pre-processing from radiance to reflectance. This variability, often lacking in single-platform or synthetic datasets, is essential to develop generalizable models capable of scaling across diverse remote sensing contexts (cherif2023spectra). While satellite-based datasets such as SpectralEarth (braham2024spectralearth) provide temporally rich but sensor-specific imagery, our dataset uniquely incorporates multi-sensor observations across spaceborne, airborne, and proximal platforms.
@@ 4 Benchmarking methods and protocols
We consider a supervised CNN-based method (cherif2023spectra; mederer2025plant) as a baseline, selected for its state-of-the-art performance in multi-trait plant prediction from a sparse annotated dataset. It is built upon EfficientNet-B0 (tan2019efficientnet) specifically framed for 1D feature extraction. The network employs multi-output regression to simultaneously predict the seven plant traits, a strategy that demonstrably outperforms single-trait modeling approaches (cherif2023spectra; scutari2014multiple).
@@
We frame the SR-GAN framework (olmschenk2019generalizing) to address hyperspectral plant trait prediction. Our implementation employs a 1D convolutional GAN architecture designed specifically for spectral data processing. In this setup, the generator learns to produce synthetic reflectance spectra, while the discriminator simultaneously performs trait regression and learns discriminative feature representations. The training objective is formulated as a composite loss that encourages the discriminator to pull real spectral samples closer in the feature space, while pushing representations of generated (synthetic) samples further apart. This contrastive learning approach allows the model to leverage unlabeled data by learning informative spectral embeddings. The overall architecture of the SR-GAN framework is illustrated in Figure 3(a) and detailed formulations of the loss components are provided in Appendix B.1.
@@
We introduce a version of the autoencoder framework proposed by (she2024spectra), which replaces the decoder with a non-learnable RTM module to reconstruct spectra, thereby integrating physical constraints into the modeling process. Specifically, our implementation employs PROSAIL-PRO (feret2021prospect), constraining the latent space to correspond directly to plant traits. PROSAIL-PRO is an RTM that combines the leaf reflectance model (PROSPECT,(jacquemoud1990prospect)) with the canopy reflectance model (4SAIL,(verhoef2007unified)). PROSPECT simulates leaf reflectance and transmittance based on biochemical composition and internal structure, while 4SAIL models the propagation of light through a vegetation canopy. Together, they simulate canopy spectral reflectance in the 400–2500 nm range using inputs such as chlorophyll content, leaf area index, and leaf angle. As previously mentioned regarding the gap between RTM-simulated and real-world spectra (§ 2), we address the inherent discrepancies between RTM-generated and observed spectra, primarily resulting from simplified geometric assumptions within the model, by implementing a learnable correction layer that refines the simulated output (she2024spectra). Our enhanced framework introduces three key improvements over the original design (she2024spectra): (1) incorporation of PROSAIL-PRO, (2) application of a supervised loss component targeting trait predictions, and (3) implementation of a composite reconstruction loss combining cosine similarity and mean absolute error to capture both spectral shape characteristics and amplitude information. The overall architecture is illustrated in Figure 3(b) and the specifications are detailed in Appendix B.2 .
@@
We adopt a MAE framework, originally designed for land cover classification (thoreau2024toulouse), to predict plant traits with hyperspectral data. The model leverages self-supervised learning by reconstructing randomly masked spectral regions, enabling the extraction of meaningful representations from unlabeled hyperspectral signatures. Similarly to the RTM-AE, our adaptation incorporates a modified reconstruction objective that combines cosine similarity and mean squared error (MSE) with appropriate weighting, allowing the model to capture both spectral shape characteristics and amplitude information. For downstream trait prediction, we attach a multi-output regression head to the latent features and fine-tune the model using labeled data. The overall architecture is illustrated in Figure 3(c) and ablation studies for the MAE architecture are provided in Appendix B.3.